how would i graph

y = 3 over 4 x + 2

Answer by checkley71(8403)

you have a slope of 3/4 (the x multiplier) & points (0,2)the y intercept of y is the right number and setting y=0 we get 0=3/4x+2 or -2=3/4x or 3x=-8 or
x=-8/3 so the second point is (-8/3,0)